# 3D Shape Generation: A Survey

### v1

@@ 6 Metrics
Evaluating the sample quality of generative models remains a challenging task [YYR+18], particularly in the context of 3D shape generation, where no single metric can reliably quantify how “good” a generated shape is [WPL+25].
@@ 6.1.1 Chamfer Distance (CD)
Popularized in 3D point cloud generation by [FSG17b], CD has since become a standard metric for evaluating geometric fidelity between a generated shape and a ground truth shape. It is used extensively in 3D reconstruction, shape completion, and point set generation, as well as in higher-level metrics such as MMD and Coverage that rely on pairwise distances.
@@ 6.1.2 Earth Mover´s Distance (EMD)
In [ADMG18], EMD was used to evaluate the quality of 3D point clouds generated by generative models by comparing the distribution of generated points to that of real data. EMD is also employed as a base distance in set-level evaluation, such as 1-NNA, MMD, and Coverage, offering a stricter measure than CD by enforcing globally optimal point-wise assignments. Consider $S_{1},S_{2}\subseteq\mathbb{R}^{3}$, the Earth Mover´s Distance is defined as:
@@ 6.1.3 F-score
where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket, returning 1 when the condition is true and 0 otherwise. $P(d)$ ranges from 0 to 100 and is interpreted as the percentage of reconstructed points that lie within a distance $d$ from the ground truth. Analogously, for each ground truth point $g\in\mathcal{G}$, its distance to the closest reconstructed point is:
@@ 6.1.4 Intersection over Union (IoU)
First applied to 3D model evaluation in voxel-based reconstructions. For instance, 3D-R2N2 [CXG+16] reported IoU between predicted and ground-truth voxels to assess single-view 3D reconstruction performance. IoU has since been ubiquitous for evaluating 3D shape prediction accuracy.
@@
$\textit{IoU}=\sum_{i,\,j,\,k}\left[I(p_{(i,\,j,\,k)}>t)\,I(y_{(i,\,j,\,k)})\right]/\sum_{i,\,j,\,k}\left[I(I(p_{(i,\,j,\,k)}>t)+I(y_{(i,\,j,\,k)}))\right]$ (17)
@@ 6.2.1 Minimum Matching Distance (MMD)
For each point cloud in the reference set, the distance to its nearest neighbor in the generated set is computed and averaged [YHH+19]:
@@ 6.2.2 Coverage (COV)
For each point cloud in the generated set, its nearest neighbor in the reference set is marked as a match [YHH+19]:
@@
According to [YHH+19], unlike COV and MMD, 1-NNA directly measures distributional similarity and takes both diversity and quality into account.
@@ 6.3.1 Fréchet PointNet Distance (FPD)
Based on the Fréchet Inception Distance (FID) proposed by [HRU+17] for images. FID measures the distance between feature distributions of real vs. generated data (using Inception-v3 features) and has been widely adopted for GAN evaluation.
[SPK19] coined the “Fréchet Point Cloud Distance”–later referred to as Fréchet PointNet Distance–by replacing 2D image features with 3D point cloud features from a PointNet classifier. tree-GAN introduced FPD as a 3D analogue of FID to evaluate the quality of generated point clouds. Subsequent works (e.g. [LWL+19, LH21a]) have used FPD, where a pretrained PointNet/PointNet++ embeds shapes and the Fréchet distance between real and generated means/covariances is computed. Lower FPD signifies closer alignment to the real shape distribution, analogous to FID.
@@ 6.3.2 CLIP-Score / FID-3D
Proposed by [HHF+21] as a reference-free metric for image captioning using OpenAI’s CLIP model. CLIP-Score computes the cosine similarity between an image and a text embedding, measuring how well a generated image matches a given description. For an image with visual CLIP embedding $v$ and a candidate caption with textual CLIP embedding $c$, $w=2.5$ is set up and computes CLIP-Score as:
[…]
Building on this idea, GET3D [GSW+22] adapted the FID to the 3D domain, introducing FID-3D, which evaluates the realism of 3D shapes by rendering them into images and computing FID over the rendered outputs. In essence, FID-3D applies the Fréchet metric to distributions of rendered images (or learned 3D features), capturing both shape and texture quality in a perceptual space.
CLIP-Score is used in text-to-3D generation to evaluate semantic alignment between a generated 3D shape (usually via its images) and the input text prompt. For example, Magic3D [LGT+23] reports CLIP-based scores to compare fidelity to the prompt, and JointDreamer [CWW+23] achieved high CLIP R-Precision (88.5%) for their text-guided 3D outputs. FID-3D was employed in GET3D to quantitatively evaluate 3D textured shape generation, and later works (e.g. [GLT+23]) use FID-3D to measure improvements in 3D shape and texture realism. Both metrics have become useful for 3D generative models that involve images or multi-modal output, connecting 3D generation to 2D image-based criteria.
@@ 7.1 Multi-Representation Learning for 3D Shape Generation
Implicit functions, for instance, cannot explicitly represent or intuitively allow editing of output surfaces as easily as meshes [SWG24]. Conversely, mesh-based representation often requires a fixed topology, particularly in approaches relying on category-specific mesh models [ZKB18, KPD19]. To address this bottleneck, methods such as  [GFK+18, WZL+18] extended mesh generation to a broader range of object categories by incorporating topological priors. While meshes are advantageous for preserving fine geometric details and enabling deformation, modifying their topology remains challenging.
Continuing along the line of 3D shape generation, works like [GSW+22] have revisited traditional “mesh and texture” representation, achieving good performance by transforming implicit surfaces to explicit meshes using techniques such as Deep Marching Tetrahedra [SGY+21].
@@ 7.3 Network Design and Large-Scale Models for 3D Shape Generation
Effective 3D shape generation requires backbone networks that can simultaneously encode complex 3D structures into latent embeddings and reconstruct detailed shapes for them [XMY23]. Developing specialized backbones has been crucial for advancing the quality and diversity of 3D shape generation. Historically, significant milestones in 3D generation have often been driven by innovation in network architectures, such as volumetric convolutional networks [ZSK+15], point-based networks like PointNet [QSMG17], and implicit function-based architectures like DeepSDF [PFS+19].
The choice of backbone architecture not only determines the underlying 3D representation but also heavily influences the generative modeling approach. For instance, voxel-based generators [WZX+16] tend to be limited by the resolution constraints, while point cloud generators [ADMG18] must address issues of unordered structure and topology.
More recently, the emergence of large-scale models has begun to impact 3D shape generation directly. Researchers are moving beyond distilling knowledge from late 2D models to optimizing 3D content and are now focusing on training native large-scale 3D generative models [LZK+24]. MeshGPT [SAA+24], for instance, adapts transformer-based architectures inspired by large language models to autoregressively generate 3D triangle meshes. However, despite the innovation, MeshGPT’s generative capabilities remain constrained by the ability and diversity of 3D training datasets, primarily producing regular objects such as furniture.
@@ 7.4 Evaluation Protocols for 3D Shape Generation
Quantitatively evaluating the quality of generated 3D shapes remains a significant and relatively unexplored problem [LZK+24]. Unlike the 2D domain, where evaluation metrics are well-established, 3D shape generation faces additional challenges due to the complexity of 3D data and the absence of standardized benchmarks.
@@ 7.5 Symmetry-aware 3D Generative Models
One of the main limitations of current generative models for 3D shapes is the lack of explicit guarantees regarding the preservation of geometric or structural properties present in the training data, such as symmetry, topological consistency, or surface continuity. Most approaches focus on learning an implicit distribution over shape space without enforcing constraints that preserve these properties [WZL+18].
@@
This leads to a methodological gap: although many 3D object categories exhibit intrinsic symmetry, generative models do not encode this property explicitly. In theory, models are expected to capture symmetry as a statistical pattern – taking into account that many objects in the training datasets are symmetrical – but there is no formal guarantee that this occurs in practice. Empirical evidence shows that generated shapes can be visually plausible yet structurally inconsistent or asymmetric.
@@
These limitations call for explicit strategies to incorporate geometric priors during generation. Symmetry, in particular, is a well-understood property in shape analysis and a natural candidate for inductive bias. While recent works have explored symmetry-aware architectures [YFZO24, KAP+25, Nag25], the area remains underdeveloped. Understanding how current models handle symmetry is thus essential for designing methods that improve structural fidelity.
@@ 8 Conclusions
In this work, we present a comprehensive survey on 3D shape generation, encompassing three core components: 3D representations, generation methods, and evaluation protocols. We begin by introducing the major categories of 3D shape representation, which define the structure and properties of the generated shapes and play a central role in model design. Next, we summarize and categorize a wide range of generation methods based on the feedforward generation paradigm. Subsequently, we summarize the datasets used to train those generation methods and the metrics used to evaluate the generation quality of the generated shapes. Finally, we discuss current challenges and outline future research directions in the field of 3D shape generation. We hope this survey offers a systematic summary of 3D shape generation, serving as a valuable reference for both researchers and newcomers in this field.

### v2

@@ 6 Metrics
Evaluating the sample quality of generative models remains a challenging task [YYR18], particularly in the context of 3D shape generation, where no single metric can reliably quantify how “good” a generated shape is [WPL25].
@@ 6.1.1 Chamfer Distance (CD)
Popularized in 3D point cloud generation by [FSG17b], CD has since become a standard metric for evaluating geometric fidelity between a generated shape and a ground truth shape. It is used extensively in 3D reconstruction, shape completion, and point set generation, as well as in higher-level metrics such as 1-NNA, MMD, and Coverage that rely on pairwise distances.
@@ 6.1.2 Earth Mover´s Distance (EMD)
In [ADMG18], EMD was used to evaluate the quality of 3D point clouds generated by generative models by comparing the distribution of generated points to that of real data. EMD is also employed as a base distance in set-level evaluation, such as 1-NNA, MMD, and Coverage, offering a stricter measure than CD by enforcing globally optimal point-wise assignments. Consider $S_{1},S_{2}\subseteq\mathbb{R}^{3}$, the Earth Mover’s Distance is defined as:
@@ 6.1.3 F-score
where $\cdot$] denotes the Iverson bracket, returning 1 when the condition is true and 0 otherwise. $P(d)$ ranges from 0 to 100 and is interpreted as the percentage of reconstructed points that lie within a distance $d$ from the ground truth. Analogously, for each ground truth point $g\in\mathcal{G}$, its distance to the closest reconstructed point is:
@@ 6.1.4 Intersection over Union (IoU)
First applied to 3D model evaluation in voxel-based reconstructions. For instance, 3D-R2N2 [CXG16] reported IoU between predicted and ground-truth voxels to assess single-view 3D reconstruction performance. IoU has since been ubiquitous for evaluating 3D shape prediction accuracy.
@@
$\textit{IoU}=\frac{\sum_{i,\,j,\,k}\left[I(p_{(i,\,j,\,k)}>t)\,I(y_{(i,\,j,\,k)})\right]}{\sum_{i,\,j,\,k}\left[I(I(p_{(i,\,j,\,k)}>t)+I(y_{(i,\,j,\,k)}))\right]}$ (17)
@@ 6.2.1 Minimum Matching Distance (MMD)
For each point cloud in the reference set, the distance to its nearest neighbor in the generated set is computed and averaged [YHH19]:
@@ 6.2.2 Coverage (COV)
For each point cloud in the generated set, its nearest neighbor in the reference set is marked as a match [YHH19]:
@@
According to [YHH19], unlike COV and MMD, 1-NNA directly measures distributional similarity and takes both diversity and quality into account.
@@ 6.2.4 Jensen-Shannon Divergence (JSD)
6.2.5 Single Shape Fréchet Inception Distance (SSFID)
Based on the Fréchet Inception Distance (FID) proposed by [HRU17] for images. FID measures the distance between feature distributions of real vs. generated data (using Inception-v3 features) and has been widely adopted for GAN evaluation.
To extend this idea to 3D, [WZ22] coined the Single Shape Fréchet Inception Distance (SSFID) for evaluating to what extent the generative model captures the patch statistics of the training shape. SSFID is defined as the FID between deep feature maps extracted from real and generated shapes.
SSIFD is commonly employed in models trained on a single 3D shape. For instance, [WZ22] trains on a voxel pyramid of the input shape and operates over a hierarchy of tri-plane feature maps, enabling the model to produce diverse shape variations across different bounding boxes and aspect ratios, while preserving large-scale structural similarity to the input shape. Similarly, Sin3DM [WLVZ23] learns the internal patch distribution from a single textured 3D shape to produce high-quality variations, and ShapeShifter [MYF25] extends diffusion models with a multiscale pipeline operating on sparse voxel grids to achieve diverse yet structurally consistent 3D generations.
@@ 6.3.1 Fréchet PointNet Distance (FPD)
Inspired by FID, tree-GAN [SPK19] introduced the “Fréchet Point Cloud Distance"–later referred to as Fréchet PointNet Distance–by replacing 2D image features with 3D point cloud features from a PointNet classifier. FPD thus serves as a 3D analogue of FID for evaluating the quality of generated point clouds. Subsequent works (e.g. [LWL19, LH21a]) have used FPD, where a pretrained PointNet/PointNet++ embeds shapes and the Fréchet distance between real and generated means/covariances is computed. Lower FPD signifies closer alignment to the real shape distribution, analogous to FID.
@@ 6.3.2 CLIP-Score / FID-3D
Proposed by [HHF21] as a reference-free metric for image captioning using OpenAI’s CLIP model. CLIP-Score computes the cosine similarity between an image and a text embedding, measuring how well a generated image matches a given description. For an image with visual CLIP embedding $v$ and a candidate caption with textual CLIP embedding $c$, $w=2.5$ is set up and computes CLIP-Score as:
$\text{CLIP-Score}(c,\,v)=w*max(cos(c,\,v),0),$ (23)
Building on this idea, GET3D [GSW22] adapted the FID to the 3D domain, introducing FID-3D, which evaluates the realism of 3D shapes by rendering them into images and computing FID over the rendered outputs. In essence, FID-3D applies the Fréchet metric to distributions of rendered images (or learned 3D features), capturing both shape and texture quality in a perceptual space.
CLIP-Score is used in text-to-3D generation to evaluate semantic alignment between a generated 3D shape (usually via its images) and the input text prompt. For example, Magic3D [LGT23] reports CLIP-based scores to compare fidelity to the prompt, and JointDreamer [CWW23] achieved high CLIP R-Precision (88.5%) for their text-guided 3D outputs. FID-3D was employed in GET3D to quantitatively evaluate 3D textured shape generation, and later works (e.g. [GLT23]) use FID-3D to measure improvements in 3D shape and texture realism. Both metrics have become useful for 3D generative models that involve images or multi-modal output, connecting 3D generation to 2D image-based criteria.
6.3.3 ULIP and Uni3D
Building on the idea of CLIP-Score, ULIP [XGX23] and Uni3D [ZWM23] have been proposed as cross-modal evaluation metrics to assess the alignment between generated 3D shapes and their corresponding textual or visual representations in a unified embedding space.
ULIP learns a joint representation of images, text, and 3D point clouds by pre-training on triplets of objects across the tree modalities. To mitigate the high demand of 3D data, ULIP leverages powerful pre-trained image and text encoders trained on massive image-text pairs, and subsequently aligns the 3D encoder to this pre-aligned image-text feature space using a small set of training triplets [XGX23]. The construction of these triplets follows a three-stage pipeline: (i) First, CAD models from ShapeNet55 [CFG15] are uniformly sampled with $N_{p}$ points, and standard point cloud data augmentation are performed. Subsequently, a 3D encoder takes the augmented point cloud $P_{i}$ as input and outputs it 3D representation $\mathbf{h^{p}_{i}}$ via $\mathbf{h^{p}_{i}}=f_{P}(P_{i})$, where $f_{P}(\cdot)$ represent the 3D backbone encoder (eg., PointNet++, PointBert and PointMLP). (ii) Since ShapeNet55 CAD models do not come with images, synthetic multi-view RGB images and depth maps are generated by placing visual cameras around each CAD model. From these, and image (or depth map) $I_{i}$ is randomly selected and encoded as $\mathbf{h^{{I}}_{i}}=f_{I}(I_{i})$, where $f_{I}(\cdot)$ is the image encoder. (iii) Finally, text descriptions are obtained from the metadata associated with each CAD model to generate textual features.
Similarly, Uni3D introduces a large multi-modal constrastive learning framework that aligns images and language with 3D point clouds, thereby providing a unified feature space suitable for both representation learning and evaluation.
Recent methods such as Hunyuan3D 2.1 [Hea25] and VolGen [TWG25] adopt ULIP and Uni3D as evaluation metrics to measure cross-modal consistency between generated 3D assets and their conditioning inputs. In particular, Hunyuan3D 2.1 reports ULIP-I and ULIP-T scores, which quatify the similarity between the point cloud and text, as well as the similarity between the point cloud and image, respectively. Likewise, Uni3D-I and Uni3D-T scores are obtained by applying the same procedure using the Uni3D model
@@ 7.1 Multi-Representation Learning for 3D Shape Generation
Implicit functions, for instance, cannot explicitly represent or intuitively allow editing of output surfaces as easily as meshes [SWG24]. Conversely, mesh-based representation often requires a fixed topology, particularly in approaches relying on category-specific mesh models [ZKB18, KPD19]. To address this bottleneck, methods such as  [GFK18, WZL18] extended mesh generation to a broader range of object categories by incorporating topological priors. While meshes are advantageous for preserving fine geometric details and enabling deformation, modifying their topology remains challenging.
Continuing along the line of 3D shape generation, works like [GSW22] have revisited traditional “mesh and texture" representation, achieving good performance by transforming implicit surfaces to explicit meshes using techniques such as Deep Marching Tetrahedra [SGY21].
@@ 7.3 Network Design and Large-Scale Models for 3D Shape Generation
Effective 3D shape generation requires backbone networks that can simultaneously encode complex 3D structures into latent embeddings and reconstruct detailed shapes for them [XMY23]. Developing specialized backbones has been crucial for advancing the quality and diversity of 3D shape generation. Historically, significant milestones in 3D generation have often been driven by innovation in network architectures, such as volumetric convolutional networks [ZSK15], point-based networks like PointNet [QSMG17], and implicit function-based architectures like DeepSDF [PFS19].
The choice of backbone architecture not only determines the underlying 3D representation but also heavily influences the generative modeling approach. For instance, voxel-based generators [WZX16] tend to be limited by the resolution constraints, while point cloud generators [ADMG18] must address issues of unordered structure and topology.
More recently, the emergence of large-scale models has begun to impact 3D shape generation directly. Researchers are moving beyond distilling knowledge from late 2D models to optimizing 3D content and are now focusing on training native large-scale 3D generative models [LZK24]. MeshGPT [SAA24], for instance, adapts transformer-based architectures inspired by large language models to autoregressively generate 3D triangle meshes. However, despite the innovation, MeshGPT’s generative capabilities remain constrained by the ability and diversity of 3D training datasets, primarily producing regular objects such as furniture.
@@ 7.4 Evaluation Protocols for 3D Shape Generation
Quantitatively evaluating the quality of generated 3D shapes remains a significant and relatively unexplored problem [LZK24]. Unlike the 2D domain, where evaluation metrics are well-established, 3D shape generation faces additional challenges due to the complexity of 3D data and the absence of standardized benchmarks.
@@ 7.5 Symmetry-aware 3D Generative Models
Figure 5: (a) Reference shapes from the ShapeNet dataset: Airplane, Car, and Chair. (b) Unconditional shape generation with 2,048 points for each class-specific model trained on PointFlow’s ShapeNet data with global normalization. Red circles highlight missing points and deformed regions.
@@
One of the main limitations of current generative models for 3D shapes is the lack of explicit guarantees regarding the preservation of geometric or structural properties present in the training data, such as symmetry, topological consistency, or surface continuity. Most approaches focus on learning an implicit distribution over shape space without enforcing constraints that preserve these properties [WZL18].
@@
This leads to a methodological gap: as can be seen in Figure 5, although many 3D objects present in widely used datasets exhibit intrinsic symmetry, generative models do not encode this property explicitly. In theory, models are expected to capture symmetry as a statistical pattern – taking into account that many objects in the training datasets are symmetrical – but there is no formal guarantee that this occurs in practice. Empirical evidence shows that generated shapes can be visually plausible yet structurally inconsistent or asymmetric.
@@
These limitations call for explicit strategies to incorporate geometric priors during generation. Symmetry, in particular, is a well-understood property in shape analysis and a natural candidate for inductive bias. While recent works have explored symmetry-aware architectures [YFZO24, KAP25, Nag25], the area remains underdeveloped. Understanding how current models handle symmetry is thus essential for designing methods that improve structural fidelity.
7.6 Efficiency and Scalability in 3D Generative Models
According to [WLVZ23], current 3D generative models rely on large-scale 3D datasets for training. However, collecting high-quality 3D data is much more expensive than images, and many artistically designed shapes have unique structures that are difficult to learn from limited collections. Moreover, state-of-the-art methods still struggle to reproduce the fine geometric details and sharp features required for digital shapes in geometry modeling [MYF25].
Table 2: Representative 3D Generative Models and their reported training/inference times. For dataset references associated with each model, please refer to table 1. For some methods, training and inference times are not explicitly reported; in those cases, we provide estimated averages derived from the model architecture, dataset scale, and GPU details.
Method GPU Details $\approx$ Training Time (GPU hours) $\approx$ Inference Time (s)
GraphCNN-GAN [VFM18] NVIDIA Quadro P6000 120 0.1
SDM-Net [GYW19] 1 $\times$ GTX 1080 Ti GPU 120 0.1
TM-Net [GWY21] 12 $\times$ RTX 2080Ti GPUs 120 0.80
Michelangelo [ZLC23] 8 $\times$ NVIDIA Tesla V100 GPUs 120 10
TAR3D [ZLL25] 8 $\times$ NVIDIA A100 GPUs 120 17.7
COD-VAE [CYJK25] 8 $\times$ RTX 4090 GPUs, 4 $\times$ A6000 GPUs 408 0.1
PointARU [MWZ25] 8 $\times$ NVIDIA Tesla V100 GPUs 125 3.21
ShapeGPT [YCZ25] 4 $\times$ NVIDIA A100 GPUs 600 0.6
PVD [ZDW21] 1 $\times$ NVIDIA Titan RTX 142 8.46
LION [ZVW22] In-house GPU cluster of NVIDIA V100 GPUs 550 27.12
SLIDE [LWA23] In-house GPU cluster of NVIDIA V100 GPUs 360 0.2
TIGER [RKLL24] In-house GPU cluster of NVIDIA V100 GPUs 164 9.73
CoPart [DDC25] 4 $\times$ NVIDIA V100 GPUs 120 65
Hunyuan3D 2.1 [Hea25] In-house GPU cluster of NVIDIA V100 GPUs 180 60
VolGen [TWG25] 16 $\times$ NVIDIA A800 GPUs 504 10
Hi3DGen [YWL25] 8 $\times$ NVIDIA A800 GPUs 500 10
SSG [WZ22] 1 $\times$ NVIDIA 3090 GPU, 1 $\times$ NVIDIA 2070 GPU 4 0.1
Sin3DGen [LCWC23] 1 $\times$ NVIDIA V100 GPU 0.25 15.8
Sin3DM [WLVZ23] 1 $\times$ NVIDIA RTX A6000 GPU 3 15.8
ShapeShifter [MYF25] 1 $\times$ NVIDIA GeForce RTX 3080 GPU 0.2 10.7
As summarized in Table 2, conventional 3D generative models (top) require substantial computational resources and involve long training and inference times. To address these limitations, a new class of 3D generative models has emerged; in fact, [WLVZ23] coined the concept of single instance generative models (SIGMs), which aim to learn the internal patch statistics from a single input instance and generate diverse new samples with similar local content.
Early attempts include DGTS [HHGC20], a GAN-based method that learns local geometric textures on mesh surfaces but not global structures, and SSG [WZ22], another GAN-based 3D shape generative model that only generates geometry from a single shape. Sin3DM [WLVZ23] further extends these approaches by adding 3 extra dimensions for RGB color and training a diffusion model on a single 3D textured shape with locally similar patterns. Sin3DGen [LCWC23] applies patch matching on the Plenoxels representation [FYT22].
More recently, ShapeShifter [MYF25] employs a multiscale diffusion framework with limited receptive fields to learn the internal structures of a given shape, building upon [KYKM23], an approach that has been used for training a generative model on a single image.
As also shown in Table 2, SIGMs (bottom) are considerably more efficient, requiring fewer computational resources and shorter training and inference times (decreasing from hundreds of GPU hours to a couple of hours or minutes on single GPUs).
We consider this a promising research direction, as it opens the possibility of high-quality 3D generation without the prohibitive costs associated with large-scale datasets and training times. By learning directly from a single example, these methods bypass the need for extensive curation pipelines, large storage infrastructures, and prolonged training and inference schedules that characterize dataset-driven approaches. Furthermore, SIGMs are particularly well-suited for domains where data is scarce, heterogeneous, or inherently unique, such as artist-designed assets, cultural heritage artifacts, or specialized industrial components.
@@ 8 Conclusions
In this work, we present a comprehensive survey on 3D shape generation, encompassing three core components: 3D representations, generation methods, and evaluation protocols. We begin by introducing the major categories of 3D shape representation, which define the structure and properties of the generated shapes and play a central role in model design. Next, we summarize and categorize a wide range of generation methods based on the feedforward generation paradigm. Subsequently, we summarize the datasets used to train those generation methods and the metrics used to evaluate the generation quality of the generated shapes. Finally, we discuss current challenges and outline future research directions in the field of 3D shape generation. We hope this survey provides a comprehensive overview of 3D shape generation, serving as a valuable reference for both researchers and newcomers in this field.
Acknowledgements
This work was supported by ANID Chile - Fondecyt Regular N° 1251263, the National Center for Artificial Intelligence CENIA FB210017, Basal ANID, and the Shape Vision Lab from the University of Chile.